# Temporal Conformal Prediction (TCP): A Distribution-Free Statistical and Machine Learning Framework for Adaptive Risk Forecasting

### v3

@@ Abstract
We propose Temporal Conformal Prediction (TCP), a distribution‐free framework for constructing well‐calibrated prediction intervals in nonstationary time series. TCP couples a modern quantile forecaster with a split‐conformal calibration layer on a rolling window and, in its TCP–RM variant, augments the conformal threshold with a single online Robbins–Monro (RM) offset to steer coverage toward a target level in real time. We benchmark TCP against GARCH, Historical Simulation, and a rolling Quantile Regression (QR) baseline across equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold). Three results are consistent across assets. First, rolling QR yields the sharpest intervals but is materially under‐calibrated (e.g., S&P 500: 86.3% vs. 95% target). Second, TCP (and TCP–RM) achieves near‐nominal coverage while being substantially narrower than Historical Simulation (e.g., S&P 500: $29\%$ width). Third, the RM update improves calibration with negligible width cost. Crisis‐window visualizations around March 2020 show TCP/TCP–RM expanding and then contracting their interval bands promptly as volatility spikes and recedes, with red dots marking days where realized returns fall outside the reported 95% interval (miscoverage). A sensitivity study confirms robustness to window size and step‐size choices. Overall, TCP provides a practical, theoretically grounded solution to calibrated uncertainty quantification under distribution shift, bridging statistical inference and machine learning for risk forecasting.
@@ 6.  RESULTS AND DISCUSSION
6.1 Main Findings: Adaptiveness and Calibration
@@
    (i) Calibration vs. sharpness trade-off. The rolling ML baseline (QR) produces the narrowest intervals across all assets but is materially under-calibrated relative to the 95% target: SP500 86.3%, BTC 89.6%, Gold 87.6%. This under-coverage is the price for sharpness (e.g., SP500 average width 2.686 versus TCP’s 3.598). In other words, rolling QR concentrates mass too aggressively in volatile regimes.
  2. 2.
    (ii) TCP/TCP-RM achieve near-nominal coverage with substantially tighter bands than Historical Simulation. TCP coverage is close to the target on all assets (SP500 94.4%, BTC 95.9%, Gold 94.3%), while keeping intervals 30–38% narrower than Historical Simulation: SP500: 3.598 vs 5.058 (-28.9%); BTC: 11.235 vs 18.058 (-37.8%); Gold: 2.663 vs 4.024 (-33.8%). Thus, TCP delivers a better calibration–sharpness balance than Hist, which attains decent coverage but with overly conservative widths.
  3. 3.
    (iii) Robbins–Monro improves calibration at essentially no cost in width. TCP-RM yields a small positive lift in coverage relative to TCP with negligible width change on all assets (e.g., SP500: +0.28 percentage points coverage, +0.009 width). This supports the utility of the online offset for fine-tuning calibration without sacrificing sharpness.
@@
GARCH under-covers consistently ($\approx$0.83–0.85) while not being as sharp as QR, suggesting its conditional-Gaussian volatility path is too slow to adapt to return asymmetries and tail thickness during stress.
@@
Crisis-window behavior. Figure 1 (SP500, Feb–Apr 2020) shows how each method behaves during the shock. The red dots mark days where the realized return falls outside the reported 95% prediction interval (i.e., miscoverage events). TCP bands widen abruptly in early March, remain elevated through the peak volatility in mid–late March, and begin to contract in April as conditions stabilize. TCP and TCP-RM deliver similar behavior and, in this short 55-day window, both achieve coverage of $\approx 0.80$ with average widths $\approx 11.55$–11.57. Rolling QR widens only modestly, keeps much tighter bands (width =7.69=7.69), and shows the most red dots around the extremes, consistent with its lower window coverage (0.709). GARCH ramps up more smoothly, reflecting its volatility recursion, and lands at the same window coverage as TCP (0.800) with narrower bands (width =9.66=9.66). Historical Simulation reacts more slowly and is the most under-covered here (0.655), despite having the narrowest bands in this window (width =5.91=5.91); this lag is expected from its long empirical lookback. Overall, the crisis snapshot illustrates a general pattern under abrupt regime change: all methods temporarily under-cover, but TCP/TCP-RM adapt fastest by inflating bands right when volatility jumps, while QR remains sharp at the cost of additional misses. This visual evidence complements the full out-of-sample results in Table 1, where TCP/TCP-RM recover to near-nominal calibration over the full evaluation horizon. Analogous crisis-window panels for BTC-USD and Gold are reported in the Appendix B.
@@
Implementation note (comparability). All learning-based methods use the same feature set. TCP/TCP-RM are evaluated out-of-sample with a rolling window of 252 points split into 192 training and 60 calibration points. Rolling QR is also evaluated out-of-sample using a 192-point trailing training window. The slight differences in the “Predictions” counts (e.g., 1509 for QR vs. 1449 for TCP/TCP-RM) reflect different warm-up requirements (QR does not reserve a calibration slice). Aligning the evaluation horizon to the longest warm-up yields qualitatively identical conclusions: QR is sharp but under-calibrated; TCP/TCP-RM are near-nominal and materially sharper than Historical Simulation.
@@ 6.2 Hyperparameter Sensitivity (SP500, TCP-RM)
100 0.005 0.9557 4.5188
100 0.010 0.9557 4.5202
100 0.050 0.9550 4.5315
252 0.005 0.9469 3.4605
252 0.010 0.9469 3.4673
252 0.050 0.9524 3.4878
500 0.005 0.9559 3.2627
500 0.010 0.9659 3.2815
500 0.050 0.9759 3.3970
What we learn. (i) Increasing $w$ consistently sharpens intervals: widths fall from $\approx 4.52$ (w=100w{=}100) to 3.263.263.403.40 (w=500w{=}500), while coverage stays at or above the target band. (ii) At a fixed $w$, larger $\gamma_{0}$ nudges coverage upward with a small width trade-off; e.g., at w=252w{=}252, $\gamma_{0}{=}0.050$ lifts coverage from $0.9469$ to $0.9524$ (+0.55pp) with a $\sim 0.03$ increase in width; at w=500w{=}500, very large $\gamma_{0}$ overshoots the target (0.9759).
Takeaway. TCP-RM is robust to reasonable choices of $(w,\gamma_{0})$. For the main experiments, we retain w=252w{=}252 and $\gamma_{0}{=}0.01$, which stay near the nominal level while keeping intervals tight; if one prefers even sharper intervals, w=500w{=}500 with $\gamma_{0}\in[0.005,0.01]$ is a good choice without material over-coverage.
@@
In Fig. 2, TCP and TCP-RM behave similarly in tranquil periods, but during stress the effective threshold $C_{\text{eff}}$ (bottom panel) rises quickly, producing wider bands (top) and steering the 30-day rolling coverage back toward the 95% target (middle). The Robbins–Monro offset $C_{\text{RM}}$ remains near zero except around regime shifts, where it provides the extra adjustment that closes the gap to target coverage. This visual evidence complements Table 1: on S&P 500, TCP-RM attains $\approx$95% coverage with only a marginal width increase relative to TCP
@@ 7.  CONCLUSION AND FUTURE WORK
Across three asset classes, TCP/TCP–RM deliver a strong calibration–sharpness balance: coverage is close to the 95% target while intervals are materially narrower than Historical Simulation. In contrast, the rolling QR baseline attains the narrowest bands but undershoots the target coverage, especially in turbulent regimes. Crisis‐window visualizations around the COVID‐19 crash illustrate TCP’s intended adaptiveness: intervals widen rapidly as volatility spikes and contract as conditions stabilize, with red dots indicating miscoverage events. A targeted sensitivity analysis shows robustness to reasonable choices of rolling window and Robbins–Monro step size, and suggests default settings that are stable in practice.
@@ 7.1 Future Work
8.  DATA AND CODE AVAILABILITY
The financial returns dataset and all code used in this paper are publicly available at: https://github.com/agnivibes/temporal-conformal-prediction-tcp.
@@ Appendix B SUPPLEMENTARY VISUALIZATIONS
Figure 3: A comparison of 95% prediction intervals from all four models for BTC-USD daily returns during the COVID-19 market crash (February-April 2020).
Figure 4: A comparison of 95% prediction intervals from all four models for Gold daily returns during the COVID-19 market crash (February-April 2020).
@@ Appendix D SUPPLEMENTARY SENSITIVITY ANALYSIS
This section provides the full sensitivity analysis results for BTC-USD (Table 3) and Gold (Table 4), complementing the main analysis on the S&P 500 presented in the results section. The findings confirm that the TCP-RM framework is robust and performs consistently across different asset classes and hyperparameter settings.
@@ Appendix E SUPPLEMENTARY SENSITIVITY ANALYSIS ON TCP (NO RM)
100 0.9538 3.7055
252 0.9517 2.6234
500 0.9417 2.3872
As with TCP-RM, larger $w$ generally yields narrower intervals. Without the adaptive offset, coverage can dip slightly below $95\%$ in some settings, underscoring the benefit of the online calibration layer.

### v4

@@ Abstract
We propose Temporal Conformal Prediction (TCP), a distribution‐free framework for constructing well‐calibrated prediction intervals in nonstationary time series. TCP couples a modern quantile forecaster with a split‐conformal calibration layer on a rolling window and, in its TCP–RM variant, augments the conformal threshold with a single online Robbins–Monro (RM) offset to steer coverage toward a target level in real time. We benchmark TCP against GARCH, Historical Simulation, and a rolling Quantile Regression (QR) baseline across equities (S&P 500), cryptocurrency (Bitcoin), and commodities (Gold). Three results are consistent across assets. First, rolling QR yields the sharpest intervals but is materially under-calibrated (e.g., S&P 500: 83.2% vs. 95% target). Second, TCP (and TCP–RM) achieves near-nominal coverage across assets, with intervals that are wider than Historical Simulation in this evaluation (e.g., S&P 500: 5.21 vs. 5.06). Third, the RM update changes calibration and width only marginally at our default hyperparameters. Crisis‐window visualizations around March 2020 show TCP/TCP–RM expanding and then contracting their interval bands promptly as volatility spikes and recedes, with red dots marking days where realized returns fall outside the reported 95% interval (miscoverage). A sensitivity study confirms robustness to window size and step‐size choices. Overall, TCP provides a practical, theoretically grounded solution to calibrated uncertainty quantification under distribution shift, bridging statistical inference and machine learning for risk forecasting.
@@ 6.  RESULTS AND DISCUSSION
Main Findings: Adaptiveness and Calibration
@@
(i) Calibration vs. sharpness trade-off. The rolling ML baseline (QR) yields the narrowest intervals across all assets but is materially under-calibrated relative to the 95% target: SP500 83.2%, BTC 85.5%, Gold 83.0%. This under-coverage is the price for sharpness (e.g., SP500 average width 3.311 versus TCP’s 5.210).
(ii) TCP/TCP–RM achieve near-nominal coverage with wider bands than Historical Simulation. TCP coverage is close to the target on all assets (SP500 95.2%, BTC 95.4%, Gold 94.3%). Relative to Historical Simulation, the intervals are wider in this setting (SP500: 5.210 vs. 5.058; BTC: 20.888 vs. 18.058; Gold: 4.737 vs. 4.024), reflecting stricter calibration under elevated volatility.
(iii) Robbins–Monro adds negligible change at these hyperparameters. TCP–RM’s coverage and width are essentially identical to TCP across assets (coverage deltas $\leq$ 0.0007; width deltas $\leq$ 0.002), indicating the online offset has minimal incremental effect here.
@@
GARCH under-covers consistently ($\approx$0.83–0.85) while being sharper than the calibrated methods, consistent with its conditional-Gaussian specification being slow to adapt to asymmetric tails during stress.
@@
Crisis-window behavior. Figure 1 (SP500, Feb–Apr 2020) shows how each method behaves during the shock. The red dots mark days where the realized return falls outside the reported 95% prediction interval (miscoverage). TCP/TCP–RM bands widen abruptly into the March spike and contract as conditions stabilize, whereas QR remains comparatively tight and under-covers around extremes. GARCH adapts smoothly via its volatility recursion, and Historical Simulation reacts with a lag due to its long lookback. Overall, all methods temporarily under-cover at the break, but TCP/TCP–RM inflate bands right when volatility jumps and recover calibration thereafter. Analogous crisis-window panels for BTC-USD and Gold are reported in Appendix B.
@@
Implementation note (comparability). All learning-based methods use the same feature set. TCP/TCP-RM are evaluated out-of-sample with a rolling window of 252 points split into 192 training and 60 calibration points. Rolling QR is also evaluated out-of-sample using a 192-point trailing training window. The slight differences in the “Predictions” counts (e.g., 1508 for QR vs. 1448 for TCP/TCP-RM) reflect different warm-up requirements (QR does not reserve a calibration slice). Aligning the evaluation horizon to the longest warm-up yields qualitatively identical conclusions: QR is sharp but under-calibrated; TCP/TCP-RM are near-nominal and, in this evaluation, wider than Historical Simulation.
@@ 6.1 Hyperparameter Sensitivity (SP500, TCP-RM)
100 0.005 0.9575 5.3052
100 0.010 0.9575 5.3066
100 0.050 0.9581 5.3172
252 0.005 0.9579 5.1168
252 0.010 0.9579 5.1191
252 0.050 0.9586 5.1314
500 0.005 0.9525 5.4135
500 0.010 0.9525 5.4319
500 0.050 0.9567 5.5348
What we learn. (i) Across these settings, w=252w{=}252 yields the narrowest intervals (width $\approx 5.12$), while w=100w{=}100 and w=500w{=}500 are wider ($\approx 5.31$ and 5.415.415.535.53, respectively), with coverage remaining near or just above the 95% target for $w\in\{100,252\}$ and slightly lower for some w=500w{=}500 configurations. (ii) At a fixed $w$, increasing $\gamma_{0}$ produces only minor changes: coverage nudges up by at most a few tenths of a percentage point with a negligible width increase (e.g., at w=252w{=}252, $\gamma_{0}{=}0.050$ lifts coverage from $0.9579$ to $0.9586$ with a $\sim 0.012$ increase in width; at w=500w{=}500, $\gamma_{0}{=}0.050$ raises coverage from $0.9525$ to $0.9567$ while widening intervals modestly).
Takeaway. TCP–RM is robust over $(w,\gamma_{0})$ in this range. For the main experiments, w=252w{=}252 with $\gamma_{0}{=}0.01$ offers a good balance (coverage $\approx 0.9579$, width $\approx 5.1191$). If one prefers higher coverage at w=500w{=}500, a slightly larger $\gamma_{0}$ (e.g., $0.05$) helps, at the cost of modestly wider bands.
@@
In Fig. 2, TCP and TCP-RM behave similarly in tranquil periods, but during stress the effective threshold $C_{\text{eff}}$ (bottom panel) rises quickly, producing wider bands (top) and steering the 30-day rolling coverage back toward the 95% target (middle). The Robbins–Monro offset $C_{\text{RM}}$ remains near zero except around regime shifts, where it provides the extra adjustment that closes the gap to target coverage. This visual evidence complements Table 1: on S&P 500, TCP-RM attains $\approx$95% coverage with only a marginal width increase relative to TCP.
@@ 7.  CONCLUSION AND FUTURE WORK
Across three asset classes, TCP/TCP–RM delivers a strong calibration–sharpness balance: coverage is close to the 95% target; in our evaluation, the intervals are wider than Historical Simulation, reflecting stricter calibration during volatile periods. In contrast, the rolling QR baseline attains the narrowest bands but undershoots the target coverage, especially in turbulent regimes. Crisis‐window visualizations around the COVID‐19 crash illustrate TCP’s intended adaptiveness: intervals widen rapidly as volatility spikes and contract as conditions stabilize, with red dots indicating miscoverage events. A targeted sensitivity analysis shows robustness to reasonable choices of rolling window and Robbins–Monro step size, and suggests default settings that are stable in practice.
@@ Appendix B SUPPLEMENTARY VISUALIZATIONS
Figure 3: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist) for BTC-USD daily returns during the COVID-19 market crash (February-April 2020).
Figure 4: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist) for Gold daily returns during the COVID-19 market crash (February-April 2020).
@@ Appendix D SUPPLEMENTARY SENSITIVITY ANALYSIS
This section provides the full sensitivity analysis results for BTC-USD (Table 3) and Gold (Table 4), complementing the main analysis on the S&P 500 presented in the results section. The findings confirm that the TCP-RM framework is robust and performs consistently across different asset classes and hyperparameter settings. For BTC-USD, widths decrease as $w$ grows (tightest at w=500w{=}500) while coverage remains near 95%–96%; $\gamma_{0}$ has negligible impact within a fixed $w$. For Gold, widths are smallest at w=100w{=}100 and increase with $w$; coverage hovers around the nominal level; $\gamma_{0}$ effects are minor at fixed $w$.
@@ Appendix E SUPPLEMENTARY SENSITIVITY ANALYSIS ON TCP (NO RM)
100 0.9487 4.6414
252 0.9454 4.7756
500 0.9492 5.0928
As with TCP-RM, the effect of $w$ is asset-dependent (e.g., BTC narrows as $w$ increases; SP500 is tightest near w=252w{=}252; Gold widens with larger $w$). Without the adaptive offset, coverage can dip slightly below 95% in some settings, underscoring the benefit of the online calibration layer.